40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 23 17:30:33 2012, jamie, nic, jenne, raji, manasa, Update, Alignment, Initial attempts to fix IFO alignment 
    Reply  Tue Oct 23 17:41:20 2012, Manasa, Update, Alignment, Power supply at OMC removed 
       Reply  Tue Oct 23 18:21:21 2012, Jenne, Update, Alignment, Power supply at OMC removed 
    Reply  Tue Oct 23 18:12:18 2012, Jenne, Update, Alignment, Tip tilt wires - the truth 
Message ID: 7603     Entry time: Tue Oct 23 18:21:21 2012     In reply to: 7600
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: Power supply at OMC removed 

Quote:

Quote:

Manasa and Raji hooked up HV power supplies to the PZTs and set them to the middle of their ranges (75 V).

 [Raji, Manasa]

The high-voltage power supply from the OMC was removed to replace one of the PZT power supplies. The power supply terminals were connected to the rear connection ports as per instructions from the manual (TB1 panel: port 3 - (-)OUT and port7 - (+)OUT). They were both switched  on and set to deliver (75V) to the PZTs.

 

 This means that the low voltage dual supply which was wired in series (so could supply a max of 63V = 2*31.5V) has been replaced with the OMC power supply.  This is okay since we haven't turned on the OMC PZTs in a long, long time.  This is *not* the power supply for the output pointing PZTs.  When she says "both", she means the new HV supply, as well as the HV supply that was already there, so both pitch and yaw for PZT2 are being supplied with 75V now.

ELOG V3.1.3-